2010-11-28 (UT)
===============

Program: CL2 (YSOs, PI: Baron et al.)

Observers: Stefan Kraus, Chris

B1  B2  B3
----------
S1  W1  E1   (W1=REF)
S2  E2  W2

(S1S2E1E2W1W2: 4-1-4-1-2-5)

CLIMB on beam 1,2,3


--- Afternoon:
Observer from last night reported oscillations on E1 and E2.
Oscillations cannot be fixed before Monday.
Therefore, we start with S1W1W2 and avoid E1 and E2.

E1 cannot be used since POP3 stucks and cannot be moved in
E1 dome drive has to be disabled due to isolation foam which was
put in by Larry in expectation of the upcoming bad weather.

Could not open the dome over the whole night
due to snow, rain, wind.